OpenStudy (anonymous):

The ratio of the length to the width of a map is 8 to 5. The map is exactly 32 inches in length. What is the width of the map?

OpenStudy (anonymous):

8/5 = 32/x (8/5)x=32 8x=160 x=20 B) 20 inches

OpenStudy (carlyleukhardt):

Nice job sunny :D

OpenStudy (anonymous):

Thank you! :)